# arXiv:2602.09029v3[cs.IT] 01 Mar 2026

### v3

@@ Abstract
We develop a sharp, experiment-level privacy theory for amplification by shuffling in the Gaussian regime: a fixed finite-output local randomizer with full support and neighboring binary datasets differing in one user. We first prove exact likelihood-ratio identities for shuffled histograms and a complete conditional-expectation linearization theorem with explicit typical-set remainders. We then derive sharp Jensen–Shannon divergence expansions, identifying the universal leading constant $I_{\pi}/(8n)$ for proportional compositions and emphasizing the correct fixed-composition covariance $\Sigma_{\pi}=(1-\pi)\Sigma_{0}+\pi\Sigma_{1}$. Next we establish equivalence to Gaussian Differential Privacy with Berry–Esseen bounds and obtain the full limiting $(\varepsilon,\delta)$ privacy curve. Finally, for unbundled multi-message shuffling we give an exact degree-$m$ likelihood ratio, asymptotic GDP formulas, and a strict bundled-versus-unbundled comparison. Further results include Local Asymptotic Normality with quantitative Le Cam equivalence, exact finite-$n$ privacy curves, a boundary Berry–Esseen theorem for randomized response, and a frequency-estimation application.
@@ 1 Introduction
The shuffle model of differential privacy interpolates between the local and central models by inserting a trusted shuffler between users and an analyst. Each of $n$ users holds a bit $x_{i}\in\{0,1\}$, applies a local randomizer $W:\{0,1\}\to\Delta(\mathcal{Y})$, and releases a message in a finite alphabet $\mathcal{Y}=\{1,\dots,d\}$. The shuffler outputs a uniformly random permutation of the $n$ messages, so the analyst observes only the permuted multiset, equivalently the histogram $N\in\mathbb{Z}_{\geq 0}^{d}$ of message counts. For a dataset with exactly $k$ ones, denote the induced histogram law by $T_{n,k}$. By exchangeability, $T_{n,k}$ depends only on $k$, and any neighboring pair of datasets induces some neighboring pair $(T_{n,k},T_{n,k+1})$.
A large literature provides amplification-by-shuffling bounds, typically at the level of rates. For protocol design, comparisons of mechanisms, and sharpness questions, one often needs constants, higher-order terms, and full privacy curves rather than order-level inequalities. This paper gives a self-contained treatment of the Gaussian regime: fixed finite-output local randomizers with full support, where neighboring shuffle experiments are asymptotically equivalent (in Le Cam distance) to a one-dimensional Gaussian shift. Two structural inputs drive the analysis: (1) exact likelihood-ratio identities that reduce shuffled comparisons to controlled functionals of a score statistic; and (2) a complete conditional-expectation linearization theorem with explicit typical-set remainders, proved via a lattice Edgeworth ratio lemma.
@@
Differential privacy was introduced by Dwork et al. [7]; see Dwork–Roth [8] for a textbook treatment. Amplification by shuffling was initiated and developed in Erlingsson et al. [10] and Cheu et al. [5], with many subsequent refinements. Feldman–McMillan–Talwar [12] provide a simple nearly optimal amplification guarantee, and Balle–Bell–Gascón–Nissim [2] introduce the privacy blanket method. Rényi divergence accounting was developed by Mironov [19] and extended to subsampled settings by Balle–Barthe–Gaboardi [1] and Wang–Balle–Kasiviswanathan [27]. Concentrated differential privacy [9, 4], composition theorems [15, 23], and numerical privacy accounting via FFT and characteristic functions [17, 29] provide complementary tools. Our analysis is complementary: rather than upper bounding leakage by general inequalities, we identify governing constants and full limiting privacy curves via likelihood-ratio and experiment-level asymptotics. The exact linear likelihood ratio identity for the canonical shuffle pair appears (in equivalent form) in Girgis et al. [13] as part of their shuffle RDP analysis. Independent and concurrent work by Takagi–Liew [24] introduces a “shuffle index” that corresponds to $\chi^{2}(W_{1}\|W_{0})$ for the canonical pair; a preliminary version of the present work appeared on Zenodo on January 3, 2026 (DOI: 10.5281/zenodo.18112202). Our framework extends to general compositions, unbundled multi-message models, and exact privacy curves.
@@ 5.2 Canonical neighboring pair: sharp $O(n^{-3})$ expansion
Finally, for the sixth absolute moment, Rosenthal’s inequality (for $p=6$; see, e.g., [20, Chapter 2, Section 9]) yields
@@ 6.3 Normal approximation under the alternative via change of measure
Then $g_{a,B}(U_{n})=U_{n}\mathbf{1}\{U_{n}\leq a\}$ almost surely and $g_{a,B}$ has bounded variation uniformly in $a$ (total variation $\leq 2B+2$). By integration by parts (Stieltjes form: $\mathbb{E}[g(X)]=g(b^{-})\mathbb{P}(X\leq b)-\int g^{\prime}(u)\mathbb{P}(X\leq u)\,du$ for $g$ of bounded variation and $X$ supported on [1,B][-1,B]; see, e.g., [11, Theorem 1, Chapter XVI]), the expectation of $g_{a,B}(U_{n})$ differs from that of $g_{a,B}(\sigma_{n}G)$ (with $G\sim N(0,1)$) by at most $\mathrm{TV}(g_{a,B})\cdot\sup_{u}|\mathbb{P}(U_{n}\leq u)-\Phi(u/\sigma_{n})|=O(n^{-1/2})$ uniformly in $a$, with the bound controlled by the Kolmogorov distance in Lemma 6.1. Since $g_{a,B}$ agrees with $u\mathbf{1}\{u\leq a\}$ on [1,B][-1,B] and $\sigma_{n}G$ exceeds $B$ only with exponentially small probability, we obtain a uniform Gaussian approximation for $\mathbb{E}[U_{n}\mathbf{1}\{U_{n}\leq a_{n}(t)\}]$ with $O(n^{-1/2})$ error. Combining the two terms yields the stated $O(n^{-1/2})$ bound; mean and variance follow by expanding $(1+u)\log(1+u)$ and $(1+u)\log^{2}(1+u)$ and using moment bounds. ∎
@@ 7.1 Trade-off and Le Cam distance for binary experiments
By the randomization theorem for binary experiments (equivalently, the Blackwell–Le Cam ordering; see, e.g., [18, Ch. 2] or [25]), these inclusions are equivalent to $\delta(E,F)\leq\eta$ and $\delta(F,E)\leq\eta$, respectively. Therefore $\Delta(E,F)\leq\eta$. ∎
@@ 9.2 Exact likelihood ratio for the canonical pair
since $L^{(m)}_{n,0}\leq w_{\max}^{m}$ pointwise by Theorem 9.1. Under $T^{(m)}_{n,0}$, $L^{(m)}_{n,0}$ is a bounded order-$m$ U-statistic in the $nm$ i.i.d. message-level ratios $w(Y_{i})$ with mean 11, so Hoeffding’s inequality for bounded U-statistics (see [21, Section 5.6]), with range upper-bounded by $w_{\max}^{m}$ (since the kernel takes values in $w_{\min}^{m},w_{\max}^{m}]$ and $w_{\max}^{m}-w_{\min}^{m}\leq w_{\max}^{m}$), gives
@@ 9.3 Canonical unbundled GDP with Berry–Esseen rate
For fixed $m$, $R^{\mathrm{deg}}_{n}$ is the sum of degenerate Hoeffding components of orders $2$ through $m$, computed over $nm$ observations. The dominant contribution (order $2$) has $\mathrm{Var}=O((nm)^{-2})$, and higher-order components have $\mathrm{Var}=O((nm)^{-j})$ for $j\geq 3$, so $\mathrm{Var}(R^{\mathrm{deg}}_{n})\leq C/(nm)^{2}$ (see [21, Chapter 5]), hence $R^{\mathrm{deg}}_{n}=O_{P}(n^{-1})$. Since $\mu_{n}(m)=\Theta(n^{-1/2})$, we have $R^{\mathrm{deg}}_{n}/\mu_{n}(m)\to 0$ in probability.
@@ A.1 A local CLT with first-order Edgeworth correction
This is a lattice Edgeworth expansion (local CLT with a polynomial correction) for sums of independent lattice-valued random vectors with bounded third moments and a uniform aperiodicity condition. In our categorical setting, all moments are bounded (each $X_{i}$ takes values in $\{e_{1},\dots,e_{d}\}$), uniform nondegeneracy $\Sigma_{\mathrm{ps}}^{-}\succ 0$ follows from (34) (every category has probability $\geq p$), and strong aperiodicity on the dual lattice holds because (34) ensures positive mass in all $d$ directions, which prevents the characteristic function from approaching $1$ on any nontrivial lattice hyperplane. See Bhattacharya–Rao [3, Theorem 20.1] and Kolassa [16, Chapter 4] for local lattice Edgeworth expansions under such conditions. ∎
@@ A.2 Ratio of local CLT evaluations
We use a local lattice Edgeworth expansion to one order higher than Lemma A.1. Because the summands are uniformly bounded (categorical) and satisfy the strong aperiodicity condition implied by (34), local Edgeworth expansions of arbitrary order hold with polynomial-growth remainders; see Bhattacharya–Rao [3, Chapter 5] and Kolassa [16, Chapter 4]. In particular, there exist polynomials $Q^{(1)}_{m}$ (cubic) and $Q^{(2)}_{m}$ (degree at most $6$), with coefficients bounded in terms of $(d,p)$, such that uniformly over $\|z\|_{2}\leq 2m^{1/8}$,

### v4

# Universal Shuffle Asymptotics: Sharp Privacy Analysis in the Gaussian Regime

@@ Abstract
We develop a sharp, experiment-level privacy theory for amplification by shuffling in the Gaussian regime: a fixed finite-output local randomizer with full support and neighboring binary datasets differing in one user. The goal is structural: to isolate the fixed full-support regime in which Gaussian/GDP is the right limit object and to make the controlling constants explicit. We first prove exact likelihood-ratio identities for shuffled histograms and a complete conditional-expectation linearization theorem with explicit typical-set remainders. We then derive sharp Jensen–Shannon divergence expansions, identifying the universal leading constant $I_{\pi}/(8n)$ for proportional compositions and emphasizing the correct fixed-composition covariance $\Sigma_{\pi}=(1-\pi)\Sigma_{0}+\pi\Sigma_{1}$. Next we establish equivalence to Gaussian Differential Privacy with Berry–Esseen bounds and obtain the full limiting $(\varepsilon,\delta)$ privacy curve. Finally, for unbundled multi-message shuffling we give an exact degree-$m$ likelihood ratio, asymptotic GDP formulas, and a strict bundled-versus-unbundled comparison. Further results include Local Asymptotic Normality with quantitative Le Cam equivalence, exact finite-$n$ privacy curves, a boundary Berry–Esseen theorem for randomized response, and a frequency-estimation application.
@@ 1 Introduction
The shuffle model of differential privacy interpolates between the local and central models by inserting a trusted shuffler between users and an analyst. Each of $n$ users holds a bit $x_{i}\in\{0,1\}$, applies a local randomizer $W:\{0,1\}\to\Delta(\mathcal{Y})$, and releases a message in a finite alphabet $\mathcal{Y}=\{1,\dots,d\}$. The neighboring input model is binary, but the output alphabet $\mathcal{Y}$ is an arbitrary fixed finite set. The shuffler outputs a uniformly random permutation of the $n$ messages, so the analyst observes only the permuted multiset, equivalently the histogram $N\in\mathbb{Z}_{\geq 0}^{d}$ of message counts. For a dataset with exactly $k$ ones, denote the induced histogram law by $T_{n,k}$. By exchangeability, $T_{n,k}$ depends only on $k$, and any neighboring pair of datasets induces some neighboring pair $(T_{n,k},T_{n,k+1})$.
A large literature provides amplification-by-shuffling bounds, and for some concrete mechanisms finite-$n$ guarantees can also be computed numerically. Those methods answer a certification question for fixed parameters: what privacy guarantee do we get for this mechanism at this $n$? They do not identify the governing regime, tell us when a Gaussian/GDP approximation is actually the right one, or explain which features of the local channel control privacy level and convergence speed. For protocol design, mechanism comparison, and sharpness questions, that structural information matters as much as the certificate itself. This paper gives a self-contained treatment of the Gaussian regime: fixed finite-output local randomizers with full support, where neighboring shuffle experiments are asymptotically equivalent (in Le Cam distance) to a one-dimensional Gaussian shift. Two structural inputs drive the analysis: (1) exact likelihood-ratio identities that reduce shuffled comparisons to controlled functionals of a score statistic; and (2) a complete conditional-expectation linearization theorem with explicit typical-set remainders, proved via a lattice Edgeworth ratio lemma.
@@
Differential privacy was introduced by Dwork et al. [8]; see Dwork–Roth [9] for a textbook treatment. Amplification by shuffling was initiated and developed in Erlingsson et al. [11] and Cheu et al. [6], with many subsequent refinements. Feldman–McMillan–Talwar [13] provide a simple nearly optimal amplification guarantee, and Balle–Bell–Gascón–Nissim [3] introduce the privacy blanket method. Numerical inversion of shuffle-amplification bounds is also used in practice; the ENPA white paper [1] is one concrete example. Rényi divergence accounting was developed by Mironov [20] and extended to subsampled settings by Balle–Barthe–Gaboardi [2] and Wang–Balle–Kasiviswanathan [28]. Concentrated differential privacy [10, 5], composition theorems [16, 24], and numerical privacy accounting via FFT and characteristic functions [18, 30] provide complementary tools. These upper-bound and numerical methods address certification for given mechanisms. Our goal here is different: to identify the fixed full-support regime itself, the correct Gaussian/GDP limit in that regime, and the constants that govern approximation and comparison. The exact linear likelihood ratio identity for the canonical shuffle pair appears (in equivalent form) in Girgis et al. [14] as part of their shuffle RDP analysis. Independent and concurrent work by Takagi–Liew [25] introduces a “shuffle index” that corresponds to $\chi^{2}(W_{1}\|W_{0})$ for the canonical pair; a preliminary version of the present work appeared on Zenodo on January 3, 2026 (DOI: 10.5281/zenodo.18112202). Our framework extends to general compositions, unbundled multi-message models, and exact privacy curves.
@@ 5.2 Canonical neighboring pair: sharp $O(n^{-3})$ expansion
Finally, for the sixth absolute moment, Rosenthal’s inequality (for $p=6$; see, e.g., [21, Chapter 2, Section 9]) yields
@@ 6.3 Normal approximation under the alternative via change of measure
Then $g_{a,B}(U_{n})=U_{n}\mathbf{1}\{U_{n}\leq a\}$ almost surely and $g_{a,B}$ has bounded variation uniformly in $a$ (total variation $\leq 2B+2$). By integration by parts (Stieltjes form: $\mathbb{E}[g(X)]=g(b^{-})\mathbb{P}(X\leq b)-\int g^{\prime}(u)\mathbb{P}(X\leq u)\,du$ for $g$ of bounded variation and $X$ supported on [1,B][-1,B]; see, e.g., [12, Theorem 1, Chapter XVI]), the expectation of $g_{a,B}(U_{n})$ differs from that of $g_{a,B}(\sigma_{n}G)$ (with $G\sim N(0,1)$) by at most $\mathrm{TV}(g_{a,B})\cdot\sup_{u}|\mathbb{P}(U_{n}\leq u)-\Phi(u/\sigma_{n})|=O(n^{-1/2})$ uniformly in $a$, with the bound controlled by the Kolmogorov distance in Lemma 6.1. Since $g_{a,B}$ agrees with $u\mathbf{1}\{u\leq a\}$ on [1,B][-1,B] and $\sigma_{n}G$ exceeds $B$ only with exponentially small probability, we obtain a uniform Gaussian approximation for $\mathbb{E}[U_{n}\mathbf{1}\{U_{n}\leq a_{n}(t)\}]$ with $O(n^{-1/2})$ error. Combining the two terms yields the stated $O(n^{-1/2})$ bound; mean and variance follow by expanding $(1+u)\log(1+u)$ and $(1+u)\log^{2}(1+u)$ and using moment bounds. ∎
@@ 7.1 Trade-off and Le Cam distance for binary experiments
By the randomization theorem for binary experiments (equivalently, the Blackwell–Le Cam ordering; see, e.g., [19, Ch. 2] or [26]), these inclusions are equivalent to $\delta(E,F)\leq\eta$ and $\delta(F,E)\leq\eta$, respectively. Therefore $\Delta(E,F)\leq\eta$. ∎
@@ 9.2 Exact likelihood ratio for the canonical pair
since $L^{(m)}_{n,0}\leq w_{\max}^{m}$ pointwise by Theorem 9.1. Under $T^{(m)}_{n,0}$, $L^{(m)}_{n,0}$ is a bounded order-$m$ U-statistic in the $nm$ i.i.d. message-level ratios $w(Y_{i})$ with mean 11, so Hoeffding’s inequality for bounded U-statistics (see [22, Section 5.6]), with range upper-bounded by $w_{\max}^{m}$ (since the kernel takes values in $w_{\min}^{m},w_{\max}^{m}]$ and $w_{\max}^{m}-w_{\min}^{m}\leq w_{\max}^{m}$), gives
@@ 9.3 Canonical unbundled GDP with Berry–Esseen rate
For fixed $m$, $R^{\mathrm{deg}}_{n}$ is the sum of degenerate Hoeffding components of orders $2$ through $m$, computed over $nm$ observations. The dominant contribution (order $2$) has $\mathrm{Var}=O((nm)^{-2})$, and higher-order components have $\mathrm{Var}=O((nm)^{-j})$ for $j\geq 3$, so $\mathrm{Var}(R^{\mathrm{deg}}_{n})\leq C/(nm)^{2}$ (see [22, Chapter 5]), hence $R^{\mathrm{deg}}_{n}=O_{P}(n^{-1})$. Since $\mu_{n}(m)=\Theta(n^{-1/2})$, we have $R^{\mathrm{deg}}_{n}/\mu_{n}(m)\to 0$ in probability.
@@ A.1 A local CLT with first-order Edgeworth correction
This is a lattice Edgeworth expansion (local CLT with a polynomial correction) for sums of independent lattice-valued random vectors with bounded third moments and a uniform aperiodicity condition. In our categorical setting, all moments are bounded (each $X_{i}$ takes values in $\{e_{1},\dots,e_{d}\}$), uniform nondegeneracy $\Sigma_{\mathrm{ps}}^{-}\succ 0$ follows from (34) (every category has probability $\geq p$), and strong aperiodicity on the dual lattice holds because (34) ensures positive mass in all $d$ directions, which prevents the characteristic function from approaching $1$ on any nontrivial lattice hyperplane. See Bhattacharya–Rao [4, Theorem 20.1] and Kolassa [17, Chapter 4] for local lattice Edgeworth expansions under such conditions. ∎
@@ A.2 Ratio of local CLT evaluations
We use a local lattice Edgeworth expansion to one order higher than Lemma A.1. Because the summands are uniformly bounded (categorical) and satisfy the strong aperiodicity condition implied by (34), local Edgeworth expansions of arbitrary order hold with polynomial-growth remainders; see Bhattacharya–Rao [4, Chapter 5] and Kolassa [17, Chapter 4]. In particular, there exist polynomials $Q^{(1)}_{m}$ (cubic) and $Q^{(2)}_{m}$ (degree at most $6$), with coefficients bounded in terms of $(d,p)$, such that uniformly over $\|z\|_{2}\leq 2m^{1/8}$,